OpenStudy (ttp):

The face of a tent is shaped like a triangle with an area of Fifteen feet squared. The area of the triangle is represented by the formula A equals one half b times h. The height of the tent is 4 feet. How wide is the base of the tent? 7.0 feet 7.5 feet 8.0 feet 8.5 feet

OpenStudy (anonymous):

do you understand where to go with the question? or do you just want an answer?

OpenStudy (ttp):

I think it's B but i'm not for sure ?

OpenStudy (ttp):

Am i correct?

OpenStudy (anonymous):

oh sorry. A= 1/2 b * h and h = 4 and A = 15. So 15/4 = 1/2 b. 15/4 * 2 = b. 15/4*2 = 15/2 = 7.5 = Answer B